# Event-Driven Safe and Resilient Control of Automated and Human-Driven Vehicles under EU-FDI Attacks

### 03 Mar 2026 v1

@@ Abstract
This paper studies the safe and resilient control of Connected and Automated Vehicles (CAVs) operating in mixed traffic environments where they must interact with Human-Driven Vehicles (HDVs) under uncertain dynamics and exponentially unbounded false data injection (EU-FDI) attacks. These attacks pose serious threats to safety-critical applications. While resilient control strategies can mitigate adversarial effects, they often overlook collision avoidance requirements. Conversely, safety-focused approaches tend to assume nominal operating conditions and lack resilience to adversarial inputs. To address these challenges, we propose a control framework that integrates event-driven Control Barrier Functions (CBFs) and Control Lyapunov Functions (CLFs) with adaptive attack-resilient control. The framework further incorporates data-driven estimation of HDV behaviors to ensure safety and resilience against EU-FDI attacks. Specifically, we focus on the lane-changing maneuver of CAVs in the presence of unpredictable HDVs and EU-FDI attacks on acceleration inputs. The event-driven approach reduces computational load while maintaining real-time safety guarantees. Simulation results, including comparisons with pure event-driven methods lacking resilience, validate the effectiveness and robustness of the proposed EDSR framework in achieving collision-free maneuvers, stable velocity regulation, and resilient operation under adversarial conditions.
@@ I INTRODUCTION
In the context of CAVs, false data injection (FDI) attacks pose a growing threat to safety-critical applications [weng2023secure]. These attacks can significantly disrupt cooperative control mechanisms by corrupting sensory or communicated data, ultimately leading to unsafe driving behaviors or even collisions. As vehicle systems increasingly rely on communication and automation, the vulnerability to such adversarial interventions becomes more pronounced.
Traditional mitigation strategies often focus on detecting and removing compromised agents to eliminate the influence of malicious data [zhang2021co, pasqualetti2013attack]. However, these approaches generally depend on restrictive assumptions, such as knowing an upper bound on the number of affected entities. In realistic and dynamic driving environments, especially in mixed traffic, such assumptions may be too strong and impractical to enforce.
@@
To address this challenge, we revisit the event-driven CBF framework in [xiao2022event] and redesign it to remain valid under EU-FDI attacks. Existing event-driven CBF methods rely on the premise that the dynamics used in barrier construction sufficiently match the true system evolution under bounded uncertainties, which guarantees meaningful Lie-derivative constraints and maintains quadratic program (QP) feasibility. This premise fundamentally breaks down under EU-FDI attacks. Exponentially unbounded adversarial inputs injected into the acceleration channel induce a rapidly growing mismatch between the modeled and actual longitudinal dynamics, which directly distorts the evolution of safety constraints. As a result, the CBF conditions may no longer certify forward invariance of the safe set, and the event-driven QP can become infeasible even though it remains valid under nominal or bounded disturbances. This distinction highlights a fundamentally different research problem from prior resilience studies. While existing EU-FDI works mainly investigate stability or consensus degradation under adversarial inputs with the control structure remaining intact, the problem addressed here concerns the loss of safety feasibility itself, where adversarial injections invalidate the safety-constrained optimization formulation. Therefore, this work does not combine existing methods but instead resolves a previously unaddressed failure mode of event-driven safety control under EU-FDI attacks. The main contributions of this paper are summarized as follows:
  • To the best of our knowledge, this paper is the first to reveal and resolve the loss of validity of event-driven CBF-based safety-constrained control under EU-FDI attacks by redesigning the safety mechanism to operate under adversarial injections. This unique combination ensures safety and resilience for CAVs interacting with HDVs in uncertain mixed traffic environments.
  • The proposed event-driven safe and resilient (EDSR) framework effectively addresses EU-FDI attacks, providing provable collision avoidance and stable operation guarantees. Compared to the pure event-driven method in [xiao2022event], simulation results demonstrate that alternative approaches fail to mitigate adversarial impacts, while the EDSR framework’s robustness and effectiveness are clearly validated under these conditions.
@@ II PROBLEM FORMULATION
In this scenario, the dynamics and control policy of the HDV are assumed to be unknown. The slow vehicle $U$ is assumed to continue moving in the slow lane at a constant speed $v_{U}$. For each CAV, indexed by $i\in{A,B}$ in Figure 1, the dynamics are defined according to the model in [he2021rule] as follows:
@@
Consider the following malicious FDI attacks targeting the acceleration control input of each CAV. Then we have
$\dot{v}_{i}(t)=\bar{u}_{i}(t)=u_{i}(t)+\gamma_{i}(t),i\in\{A,B\}$ (2)
@@
$x(t)\in\mathbb{R}$ is uniformly ultimately bounded (UUB) with the ultimate bound $b$, if there exist constants $b,c>0$, independent of $t_{0}\geq 0$, and for every $a\in(0,c)$, there exists $t_{1}=t_{1}(a,b)\geq 0$, independent of $t_{0}$, such that $\left|x\left(t_{0}\right)\right|\leq c\Rightarrow|x(t)|\leq b,\quad\forall t\geq t_{0}+t_{1}.$
Now, we introduce the event-driven safe and resilient control problem.
@@
where $v_{d}(t)$ is the desired cruising speed in the fast lane, and $l$ is the centerline of the target lane. $\alpha_{u}$, $\alpha_{l}$, and $\alpha_{v}$ are non-negative weights that represent the trade-offs between energy consumption, lateral deviation from the target lane, and velocity tracking performance, respectively.
@@ III HUMAN-IN-THE-LOOP SAFE LANE-CHANGING MANEUVER
In this section, an attack-resilient event-driven safe control framework is proposed, which incorporates an attack-resilient controller augmented with event-driven CBFs and CLFs, to address the EDSR control problem defined in Problem 1. To enforce safety under adversarial conditions, the original state constraints in the EDSR control problem are replaced by CBF-based constraints. Specifically, for any state constraint of the form $b(x)\geq 0$, a corresponding CBF condition is introduced.
@@
Solving the EDSR control problem 1 with event-driven CBFs requires accurately estimating the unknown dynamics of HDVs. With these HDV state estimates, the CAVs then enforce safety by transforming original constraints into corresponding CBF constraints at every control update. By doing so, the original EDSR control problem can be recast into a series of QPs, similar in form to those discussed in [ames2014control]. An event-driven method further enhances efficiency by adaptively determining the appropriate triggering moments for solving these QPs.
@@ III-A Event-Driven Safety Control
To ensure that the constraints in the EDSR control problem—namely, the speed and control constraints (3), position constraints (4), and safety constraints (5)—are satisfied at all times, CBFs are employed. Each of these constraints can be expressed as $b_{n}(\boldsymbol{x}(t))\geq 0$ for $n\in\{1,2,\ldots,N\}$, where $\boldsymbol{x}(t)$ includes all relevant vehicle states, i.e., $\boldsymbol{x}(t)=\{\boldsymbol{x}_{A}(t),\boldsymbol{x}_{B}(t),\boldsymbol{x}_{H}(t),\boldsymbol{x}_{U}(t)\}$. The corresponding CBF condition for each constraint takes the general form:
@@
$V_{1}(\boldsymbol{x}_{B}(t))=(v_{B}(t)-v_{d}(t))^{2},$
$V_{2}(\boldsymbol{x}_{A}(t))=(v_{A}(t)-v_{d}(t))^{2},$
$V_{3}(\boldsymbol{x}_{B}(t))=(y_{B}(t)-l)^{2},V_{4}(\boldsymbol{x}_{A}(t))=(y_{A}(t)-l)^{2},$
@@
where the first inequality corresponds to the longitudinal channel (acceleration control), and the second corresponds to the lateral channel (steering control), with $i\in\{A,B\}$ and $j\in\{1,2,3,4\}$. The variables $\delta_{j}$ are relaxation terms introduced to preserve QP feasibility.
@@
$\begin{aligned} &t_{k+1}=\min\left\{t>t_{k}:|\boldsymbol{e}(t)|=\boldsymbol{w},\text{ or }|\dot{\boldsymbol{e}}(t)|=\boldsymbol{\nu},\right.\\ &\text{ or }\left|\boldsymbol{x}_{i}(t)-\boldsymbol{x}_{i}\left(t_{k}\right)\right|=\boldsymbol{s}_{i},\text{ or }\left.\left|\overline{\boldsymbol{x}}_{H}(t)-\overline{\boldsymbol{x}}_{H}\left(t_{k}\right)\right|=\boldsymbol{s}_{H}\right\},\end{aligned}$ (18)
@@
Although the triggering condition in (18) appears mathematically involved, all terms are directly measurable from onboard sensors, including the state estimation error, its derivative, and the state deviation from the last update instant. No additional optimization or model-based computation is required to evaluate the triggering rule. Therefore, the event-driven mechanism is computationally lightweight and fully implementable in real time. In implementation, time is discretized with a fixed sampling period $T_{s}$, and the control problem is solved at time instants $t_{k}=t_{0}+kT_{s}$, for $k=1,2,\ldots$. At each update instant, the EDSR control problem 1 is reformulated as the QP (17). This optimization ensures that the control input not only respects all safety constraints but also minimizes the deviation from the nominal control input. The solution is obtained by solving the QP in [zhang2023novel].
@@ III-B Attack-Resilient Control
Based on Lemma 1 and Lemma 2, the CAV achieves optimal trajectory tracking by minimizing control energy and terminal deviation from the desired lane position and speed, while the state $x_{i}(t)$ remains within a time-varying elliptical safe region for all $t\geq 0$. This safety is enforced through a QP that integrates CBFs and CLFs under an event-driven framework. Hence, the first two objectives of the EDSR control problem are achieved. Next, we prove that the speed regulation error $\varepsilon_{i}(t)$ for each CAV is UUB. A Lyapunov function candidate for each CAV $i\in\{A,B\}$ is defined as $V(\varepsilon_{i}(t))=(v_{i}(t)-v_{d}(t))^{2}$. Then, one has
$\dot{V}(\varepsilon_{i}(t))=$ $2(v_{i}(t)-v_{d}(t))\dot{v}_{i}(t)$ (23)
@@
By using the integral solution for differential equations in (22), one has $\exp\left(\hat{\rho}_{i}(t)\right)=\exp\left(\hat{\rho}_{i}(0)\right)\exp\left(\alpha_{i}\int_{0}^{t}\left|\varepsilon_{i}(\tau)\right|\mathrm{d}\tau\right)$. Define the compact set $\Upsilon_{i}\triangleq\{\left|\varepsilon_{i}(t)\right|\leq\kappa_{i}\}$. Pick $\hat{\rho}_{i}(0)=0$ so that $\exp\left(\hat{\rho}_{i}(0)\right)=1$. Then, outside the compact set $\Upsilon_{i}$, one has that there exists a time $t_{1}$ such that for all $t\geq t_{1}$, $|\varepsilon_{i}(t)||\gamma_{i}(t)|-|\varepsilon_{i}(t)|\exp(\hat{\rho}_{i}(t))\leq 0$. Considering Assumption 1, one has $\lim_{t\to\infty}\exp(-c_{i}t^{2})|\gamma_{i}(t)|=0$. Hence, one obtains that outside the compact set $\Upsilon_{i}$, for all $t\geq t_{1}$, $\dot{V}(\varepsilon_{i}(t))\leq 0$. Then, by LaSalle’s invariance principle [lasalle1960some], one has that $\varepsilon_{i}(t)$ is UUB. This completes the proof. ∎
@@ IV SIMULATION RESULTS
This section presents simulation results that demonstrate the effectiveness of the proposed framework in enabling each CAV to perform optimal lane-changing maneuvers with safety guarantees, even in the presence of HDVs whose dynamics are unknown to the CAVs. The selected lane-changing scenario is intentionally designed as a worst-case mixed-traffic interaction, where safety constraints, human uncertainty, and adversarial actuator corruption coexist. Unlike many existing works that assume cooperative or predictable surrounding vehicles, the considered scenario includes an HDV whose behavior is unknown and cannot be influenced by the controller, thereby introducing significant uncertainty into the traffic environment. Under EU-FDI attacks that corrupt the longitudinal dynamics of the CAVs, this unpredictability prevents reliance on structured interactions or cooperative maneuvers for safety recovery. Consequently, the scenario forms a highly adversarial test condition that highlights the necessity of the proposed EDSR framework in realistic mixed-traffic environments. The simulation setting corresponds to the mixed-traffic lane-changing scenario shown in Figure 1 and Figure 2. Vehicle $U$ is assumed to travel at a constant speed of $v_{U}(t)=20~\mathrm{m/s}$ throughout the simulation, although this assumption is not required by the proposed control framework. The dynamics of vehicles $A$ and $B$ follow the model described in (1). The allowable speed range for the CAVs is specified as $v(t)\in[15,35]~\mathrm{m/s}$, and their acceleration and steering inputs are constrained within $\boldsymbol{\xi}(t)\in[(-7,-\pi/4),(3.3,\pi/4)]~\mathrm{m/s}^{2}$. The desired speed $v_{d}(t)$ is set to $30~\mathrm{m/s}$ to reflect the traffic flow velocity. To ensure safety, elliptical safe regions defined in (5) are employed, where the ellipse parameters are set as $a_{A}=a_{B}=0.6$ to represent the reaction time of the CAVs, and $b_{A}=b_{B}=0.1$ to approximate the lane width $l=4~\mathrm{m}$ via the minor axis. The maximum allowable time for completing the lane-changing maneuver is set to $T_{f}=15~\mathrm{s}$. The real dynamics of the HDV $H$, which are unknown to the controller, are given by
@@
where $u_{H}(t)$ and $\phi_{H}(t)$ represent either a random policy or manual control by a human driver input, and the disturbances $\varepsilon_{1}(t)\in[-0.7,0.7]$, $\varepsilon_{2}(t)\in[-0.5,0.5]$, $\varepsilon_{3}(t)\in[-0.5,0.5]$, and $\varepsilon_{4}(t)\in[-0.7,0.7]$ are modeled as uniformly distributed random variables. The initial states of the vehicles at time $t_{0}=0$ are given as $\boldsymbol{x}_{A}(t_{0})=[50~\mathrm{m},4~\mathrm{m},0~\mathrm{rad},29~\mathrm{m/s}]^{T}$, $\boldsymbol{x}_{B}(t_{0})=[20~\mathrm{m},0~\mathrm{m},0~\mathrm{rad},25~\mathrm{m/s}]^{T}$, $\boldsymbol{x}_{H}(t_{0})=[10~\mathrm{m},4~\mathrm{m},0~\mathrm{rad},28~\mathrm{m/s}]^{T}$, and $\boldsymbol{x}_{U}(t_{0})=[60~\mathrm{m},0~\mathrm{m},0~\mathrm{rad},20~\mathrm{m/s}]^{T}$. The weights in the objective function (17) are selected as $\alpha_{u_{A}^{s}}=\alpha_{u_{B}^{s}}=1$, $p_{1}=p_{2}=p_{4}=1$, and $p_{3}=100$. The initial state of the adaptive HDV model (12) is set to match the true HDV state, i.e., $\overline{\boldsymbol{x}}_{H}(t_{0})=\boldsymbol{x}_{H}(t_{0})$, and the adaptive terms are initialized as $h_{x}(t_{0})=h_{y}(t_{0})=h_{\theta}(t_{0})=h_{v}(t_{0})=0$. The parameters $s_{i}$ in Lemma 2 are set as $s_{i}=[0.01\,\mathrm{m},0.005\,\mathrm{m},0.01\,\mathrm{rad},1\,\mathrm{m/s}]$ for all vehicles $i\in\{A,B,H,U\}$. The corresponding bounds for the state error $\boldsymbol{e}(t)$ and its derivative $\dot{\boldsymbol{e}}(t)$ are $\boldsymbol{w}=[0.2\,\mathrm{m},0.1\,\mathrm{m},0.1\,\mathrm{rad},1\,\mathrm{m/s}]$ and $\boldsymbol{\nu}=[0.5\,\mathrm{m/s},0.2\,\mathrm{m/s},0.1\,\mathrm{rad/s},1\,\mathrm{m/s^{2}}]$, respectively. Additionally, the allowable error threshold to terminate the maneuver is set as $\sigma=0.3\,\mathrm{m}$.

### 09 Mar 2026 v2

@@ Abstract
This paper studies the safe and resilient control of Connected and Automated Vehicles (CAVs) operating in mixed traffic environments where they must interact with Human-Driven Vehicles (HDVs) under uncertain dynamics and exponentially unbounded false data injection (EU-FDI) attacks. These attacks pose serious threats to safety-critical applications. While resilient control strategies can mitigate adversarial effects, they often overlook collision avoidance requirements. Conversely, safety-critical approaches tend to assume nominal operating conditions and lack resilience to adversarial inputs. To address these challenges, we propose an event-driven safe and resilient (EDSR) control framework that integrates event-driven Control Barrier Functions (CBFs) and Control Lyapunov Functions (CLFs) with adaptive attack-resilient control. The framework further incorporates data-driven estimation of HDV behaviors to ensure safety and resilience against EU-FDI attacks. Specifically, we focus on the lane-changing maneuver of CAVs in the presence of unpredictable HDVs and EU-FDI attacks on acceleration inputs. The event-driven approach reduces computational load while maintaining real-time safety guarantees. Simulation results, including comparisons with conventional safety-critical control methods that lack resilience, validate the effectiveness and robustness of the proposed EDSR framework in achieving collision-free maneuvers, stable velocity regulation, and resilient operation under adversarial conditions.
@@ I INTRODUCTION
In the context of CAVs, false data injection (FDI) attacks pose a growing threat to safety-critical applications [weng2023secure]. These attacks can significantly disrupt cooperative control mechanisms by corrupting sensory or communicated data, ultimately leading to unsafe driving behaviors or even collisions. As vehicle systems increasingly rely on communication and automation, the vulnerability to such adversarial interventions becomes more pronounced. Traditional mitigation strategies often focus on detecting and removing compromised agents to eliminate the influence of malicious data [zhang2021co, pasqualetti2013attack]. However, these approaches generally depend on restrictive assumptions, such as knowing an upper bound on the number of affected entities. In realistic and dynamic driving environments, especially in mixed traffic, such assumptions may be too strong and impractical to enforce.
@@
To address this challenge, we revisit the event-driven CBF framework in [xiao2022event] and redesign it to remain valid under EU-FDI attacks. Existing event-driven CBF methods rely on the premise that the dynamics used in barrier construction sufficiently match the true system evolution under bounded uncertainties, which guarantees meaningful Lie-derivative constraints and maintains quadratic program (QP) feasibility. This premise fundamentally breaks down under EU-FDI attacks. Exponentially unbounded adversarial inputs injected into the acceleration channel induce a rapidly growing mismatch between the modeled and actual longitudinal dynamics, which directly distorts the evolution of safety constraints. As a result, the CBF conditions may no longer certify forward invariance of the safe set, and the event-driven QP can become infeasible even though it remains valid under nominal or bounded disturbances. This distinction highlights a fundamentally different research problem from prior resilience studies. While existing EU-FDI works mainly investigate stability or consensus degradation under adversarial inputs with the control structure remaining intact, the problem addressed here concerns the loss of safety feasibility itself, where adversarial injections invalidate the safety-constrained optimization formulation. Therefore, this work does not combine existing methods but instead resolves a previously unaddressed failure mode of event-driven safety-critical control under EU-FDI attacks. The main contributions of this paper are summarized as follows:
  • To the best of our knowledge, this paper is the first to reveal and resolve the loss of validity of event-driven CBF-based safety-critical control under EU-FDI attacks by redesigning the safety mechanism to operate under adversarial injections. This unique combination ensures safety and resilience for CAVs interacting with HDVs in uncertain mixed traffic environments.
  • The proposed EDSR framework effectively addresses EU-FDI attacks, providing provable collision avoidance and stable operation guarantees. Compared to the conventional safety-critical control method in [xiao2022event], simulation results demonstrate that alternative approaches fail to mitigate adversarial impacts, while the EDSR framework’s robustness and effectiveness are clearly validated under these conditions.
@@ II PROBLEM FORMULATION
In this scenario, the dynamics and control policy of the HDV are assumed to be unknown. The slow vehicle $U$ is assumed to continue moving in the slow lane at a constant speed $v_{U}$. For each CAV, indexed by $i\in\{A,B\}$ in Figure 1, the dynamics are defined according to the model in [he2021rule] as follows:
@@
Consider the following malicious FDI attacks targeting the acceleration control input of each CAV:
$\dot{v}_{i}(t)=\bar{u}_{i}(t)=u_{i}(t)+\gamma_{i}(t),\;i\in\{A,B\},$ (2)
@@
$x(t)\in\mathbb{R}$ is uniformly ultimately bounded (UUB) with the ultimate bound $b$, if there exist constants $b,c>0$, independent of $t_{0}\geq 0$, and for every $a\in(0,c)$, there exists $t_{1}=t_{1}(a,b)\geq 0$, independent of $t_{0}$, such that if $\left|x\left(t_{0}\right)\right|\leq a$, then $|x(t)|\leq b,\quad\forall t\geq t_{0}+t_{1}.$
Now, we introduce the safe and attack-resilient (SAR) control problem.
@@
where $v_{d}$ is the desired cruising speed in the fast lane, and $l$ is the centerline of the target lane. $\alpha_{u}$, $\alpha_{l}$, and $\alpha_{v}$ are non-negative weights that represent the trade-offs between energy consumption, lateral deviation from the target lane, and velocity tracking performance, respectively.
@@ III HUMAN-IN-THE-LOOP SAFE LANE-CHANGING MANEUVER
In this section, an attack-resilient event-driven safe control framework is proposed, which incorporates an attack-resilient controller augmented with event-driven CBFs and CLFs, to address the SAR control problem defined in Problem 1. To enforce safety under adversarial conditions, the original state constraints in Problem 1 are replaced by CBF-based constraints. Specifically, for any state constraint of the form $b(x)\geq 0$, a corresponding CBF condition is introduced.
@@
Solving the SAR control problem 1 with event-driven CBFs requires accurately estimating the unknown dynamics of HDVs. With these HDV state estimates, the CAVs then enforce safety by transforming original constraints into corresponding CBF constraints at every control update. By doing so, the original SAR control problem can be recast into a series of QPs, similar in form to those discussed in [ames2014control]. An event-driven method further enhances efficiency by adaptively determining the appropriate triggering moments for solving these QPs.
@@ III-A Event-Driven Safety Control
To ensure that the constraints in the SAR control problem—namely, the speed and control constraints (3), position constraints (4), and safety constraints (5)—are satisfied at all times, CBFs are employed. Each of these constraints can be expressed as $b_{n}(\boldsymbol{x}(t))\geq 0$ for $n\in\{1,2,\ldots,N\}$, where $\boldsymbol{x}(t)$ includes all relevant vehicle states, i.e., $\boldsymbol{x}(t)=\{\boldsymbol{x}_{A}(t),\boldsymbol{x}_{B}(t),\boldsymbol{x}_{H}(t),\boldsymbol{x}_{U}(t)\}$. The corresponding CBF condition for each constraint takes the general form:
@@
$V_{1}(\boldsymbol{x}_{B}(t))=(v_{B}(t)-v_{d})^{2},V_{2}(\boldsymbol{x}_{A}(t))=(v_{A}(t)-v_{d})^{2},$
$V_{3}(\boldsymbol{x}_{B}(t))=(y_{B}(t)-l)^{2},V_{4}(\boldsymbol{x}_{A}(t))=(y_{A}(t)-l)^{2}.$
@@
where the first inequality corresponds to the longitudinal channel (acceleration control), and the second corresponds to the lateral channel (steering control), with $i\in\{A,B\}$ and $j\in\{1,2,3,4\}$. Here, $u_{i}^{s}(t)$ denotes the nominal safety control input and the variables $\delta_{j}$ are relaxation terms introduced to preserve QP feasibility.
@@
$\begin{aligned} &t_{k+1}=\min\{t>t_{k}:\;\exists j\in\{x,y,\theta,v\}\ \text{s.t.}\ |e_{j}(t)|\geq w_{j},\\ &\text{or }\exists j\in\{x,y,\theta,v\}\ \text{s.t.}\ |\dot{e}_{j}(t)|\geq\nu_{j},\\ &\text{or }|\boldsymbol{x}_{i}(t)-\boldsymbol{x}_{i}(t_{k})|\geq\boldsymbol{s}_{i},\text{or }|\overline{\boldsymbol{x}}_{H}(t)-\overline{\boldsymbol{x}}_{H}(t_{k})|\geq\boldsymbol{s}_{H}\},\end{aligned}$ (18)
@@
Although the triggering condition in (18) appears mathematically involved, all terms are directly measurable from onboard sensors, including the state estimation error, its derivative, and the state deviation from the last update instant. No additional optimization or model-based computation is required to evaluate the triggering rule. Therefore, the event-driven mechanism is computationally lightweight and fully implementable in real time. In implementation, time is discretized with a fixed sampling period $T_{s}$, and the control problem is solved at time instants $t_{k}=t_{0}+kT_{s}$, for $k=1,2,\ldots$. At each update instant, the SAR control problem 1 is reformulated as the QP (17). This optimization ensures that the control input not only respects all safety constraints but also minimizes the deviation from the nominal control input. The solution is obtained by solving the QP in [zhang2023novel].
@@ III-B Attack-Resilient Control
Based on Lemma 1 and Lemma 2, the CAV achieves optimal trajectory tracking by minimizing control energy and terminal deviation from the desired lane position and speed, while the state $x_{i}(t)$ remains within a time-varying elliptical safe region for all $t\geq 0$. This safety is enforced through a QP that integrates CBFs and CLFs under an event-driven framework. Hence, the first two objectives of the SAR control problem are achieved. Next, we prove that the speed regulation error $\varepsilon_{i}(t)$ for each CAV is UUB. A Lyapunov function candidate for each CAV $i\in\{A,B\}$ is defined as $V(\varepsilon_{i}(t))=(v_{i}(t)-v_{d})^{2}$. Then, one has
$\dot{V}(\varepsilon_{i}(t))=$ $2(v_{i}(t)-v_{d})\dot{v}_{i}(t)$ (23)
@@
By using the integral solution for differential equations in (22), one has $\exp\left(\hat{\rho}_{i}(t)\right)=\exp\left(\hat{\rho}_{i}(0)\right)\exp\left(\alpha_{i}\int_{0}^{\top}\left|\varepsilon_{i}(\tau)\right|\mathrm{d}\tau\right)$. Define the compact set $\Upsilon_{i}\triangleq\{\left|\varepsilon_{i}(t)\right|\leq\kappa_{i}\}$. Pick $\hat{\rho}_{i}(0)=0$ so that $\exp\left(\hat{\rho}_{i}(0)\right)=1$. Then, outside the compact set $\Upsilon_{i}$, one has that there exists a time $t_{1}$ such that for all $t\geq t_{1}$, $|\varepsilon_{i}(t)||\gamma_{i}(t)|-|\varepsilon_{i}(t)|\exp(\hat{\rho}_{i}(t))\leq 0$. Considering Assumption 1, one has $\lim_{t\to\infty}\exp(-c_{i}t^{2})|\gamma_{i}(t)|=0$. Hence, one obtains that outside the compact set $\Upsilon_{i}$, for all $t\geq t_{1}$, $\dot{V}(\varepsilon_{i}(t))\leq 0$. Then, by LaSalle’s invariance principle [lasalle1960some], one has that $\varepsilon_{i}(t)$ is UUB. This completes the proof. ∎
@@ IV SIMULATION RESULTS
This section presents simulation results that demonstrate the effectiveness of the proposed framework in enabling each CAV to perform optimal lane-changing maneuvers with safety guarantees under EU-FDI attacks, even in the presence of HDVs whose dynamics are unknown to the CAVs. The selected lane-changing scenario is intentionally designed as a worst-case mixed-traffic interaction, where safety constraints, human uncertainty, and adversarial actuator corruption coexist. Unlike many existing works that assume cooperative or predictable surrounding vehicles, the considered scenario includes an HDV whose behavior is unknown and cannot be influenced by the controller, thereby introducing significant uncertainty into the traffic environment. Under EU-FDI attacks that corrupt the longitudinal dynamics of the CAVs, this unpredictability prevents reliance on structured interactions or cooperative maneuvers for safety recovery. Consequently, the scenario forms a highly adversarial test condition that highlights the necessity of the proposed EDSR framework in realistic mixed-traffic environments. The simulation setting corresponds to the mixed-traffic lane-changing scenario shown in Figure 1 and Figure 2. Vehicle $U$ is assumed to travel at a constant speed of $v_{U}(t)=20~\mathrm{m/s}$ throughout the simulation, although this assumption is not required by the proposed control framework. The dynamics of vehicles $A$ and $B$ follow the model described in (1). The allowable speed range for the CAVs is specified as $v_{i}(t)\in[15,35]~\mathrm{m/s}$, and their acceleration and steering inputs are constrained within $\boldsymbol{\xi}_{i}(t)\in[(-7,-\pi/4),(3.3,\pi/4)]~\mathrm{m/s}^{2}$. The desired speed $v_{d}$ is set to $30~\mathrm{m/s}$ to reflect the traffic flow velocity. To ensure safety, elliptical safe regions defined in (5) are employed, where the ellipse parameters are set as $a_{A}=a_{B}=0.6$ to represent the reaction time of the CAVs, and $b_{A}=b_{B}=0.1$ to approximate the lane width $l=4~\mathrm{m}$ via the minor axis. The maximum allowable time for completing the lane-changing maneuver is set to $T_{f}=15~\mathrm{s}$. The real dynamics of the HDV $H$, which are unknown to the controller, are given by
@@
where $u_{H}(t)$ and $\phi_{H}(t)$ represent either a random policy or manual control by a human driver input, and the disturbances $\varepsilon_{1}(t)\in[-0.7,0.7]$, $\varepsilon_{2}(t)\in[-0.5,0.5]$, $\varepsilon_{3}(t)\in[-0.5,0.5]$, and $\varepsilon_{4}(t)\in[-0.7,0.7]$ are modeled as uniformly distributed random variables. The initial states of the vehicles at time $t_{0}=0$ are given as $\boldsymbol{x}_{A}(t_{0})=[50~\mathrm{m},4~\mathrm{m},0~\mathrm{rad},29~\mathrm{m/s}]^{\top}$, $\boldsymbol{x}_{B}(t_{0})=[20~\mathrm{m},0~\mathrm{m},0~\mathrm{rad},25~\mathrm{m/s}]^{\top}$, $\boldsymbol{x}_{H}(t_{0})=[10~\mathrm{m},4~\mathrm{m},0~\mathrm{rad},28~\mathrm{m/s}]^{\top}$, and $\boldsymbol{x}_{U}(t_{0})=[60~\mathrm{m},0~\mathrm{m},0~\mathrm{rad},20~\mathrm{m/s}]^{\top}$. The weights in the objective function (17) are selected as $\alpha_{u_{A}^{s}}=\alpha_{u_{B}^{s}}=1$, $p_{1}=p_{2}=p_{4}=1$, and $p_{3}=100$. The initial state of the adaptive HDV model (12) is set to match the true HDV state, i.e., $\overline{\boldsymbol{x}}_{H}(t_{0})=\boldsymbol{x}_{H}(t_{0})$, and the adaptive terms are initialized as $h_{x}(t_{0})=h_{y}(t_{0})=h_{\theta}(t_{0})=h_{v}(t_{0})=0$. The parameters $s_{i}$ in Lemma 2 are set as $s_{i}=[0.01\,\mathrm{m},0.005\,\mathrm{m},0.01\,\mathrm{rad},1\,\mathrm{m/s}]^{\top}$ for all vehicles $i\in\{A,B,H,U\}$. The corresponding bounds for the state error $\boldsymbol{e}(t)$ and its derivative $\dot{\boldsymbol{e}}(t)$ are $\boldsymbol{w}=[0.2\,\mathrm{m},0.1\,\mathrm{m},0.1\,\mathrm{rad},1\,\mathrm{m/s}]^{\top}$ and $\boldsymbol{\nu}=[0.5\,\mathrm{m/s},0.2\,\mathrm{m/s},0.1\,\mathrm{rad/s},1\,\mathrm{m/s^{2}}]^{\top}$, respectively. Additionally, the allowable error threshold to terminate the maneuver is set as $\sigma=0.3\,\mathrm{m}$.